# Fairness and Bias in Multimodal : A Survey

### v1

@@ 4.2 LLM s
The introduction of the Generative Pretrained Transformer (GPT-2) Radford et al. (2019) was a turning point in the LLM lanscape with its 1.5B parameters. As pointed out earlier, training such a large model required a lot of data and the Internet-sourced WebText was used for this purpose. Updated versions of the dataset have also been used for its recent successors, including GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023). The attendant problems of bias witnessed in GPT-2 Gehman et al. (2020) have followed successive versions.
Quantitative bias measurement in NLP may be placed into 3 categories: preuse, as carried out by Alkhaled et al. (2023) and Adewumi et al. (2023b), intrinsic Caliskan et al. (2017); May et al. (2019), and extrinsic, as observed by Ramesh et al. (2023). The first, second and third involve quantifying bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively. Far more work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al. (2023). For example, Delobelle et al. (2022) and Welbl et al. (2021) measured bias in pre-trained language models. Others have surveyed fairness and bias in LLM s Bender et al. (2021); Meade et al. (2022); Gallegos et al. (2023); Chang et al. (2024); Myers et al. (2024).
@@
In other works, Schramowski et al. (2022) show that moral directions, i.e. what is morally right or wrong to do, are present in LLM s. It is, however, highly debatable if models can be considered moral agents. Also, Santurkar et al. (2023) explored whose opinions LLM s reflect. With regards to healthcare and bias, work has also been done Harrer (2023); Nashwan and Abujaber (2023). In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias.
Bias in multilingual AI:
@@
Some of the multimodal data in Table 2 contain multilingual data. Such result in multilingual models and embeddings. For example, CC3M-35L, COCO-35L, and WebLI, used to train PaLI. WebLI is a mix of pre-training tasks with texts in 109 languages Chen et al. (2022). Some of these languages fall in the category of low-resource languages Adewumi et al. (2023a). Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words, using WEAT Caliskan et al. (2017). In the study by Wambsganss et al. (2022), they found that the pretrained German language models, GermanBERT, GermanT5, and German GPT-2, had substantial conceptual, racial, and gender bias. This was also confirmed by Kraft et al. (2022), who observed sexist stereotypes in some of the models (e.g. family- and care-related terms related to female while crime and perpetrators related to male). Similarly, for Dutch, Delobelle et al. (2020) investigated gender and occupation biases in RobBERT (a Dutch RoBERTa Liu et al. (2019)), through a template-based association test Kurita et al. (2019); May et al. (2019). Huang et al. (2020) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
@@ 5 Discussion
Although there are many limitations or risks of multimodal AI or LLM s Acosta et al. (2022); Adewumi et al. (2024b); Pettersson et al. (2024); Adewumi et al. (2024a), perhaps, the issue of fairness and bias rank among the topmost Mehrabi et al. (2021). In addition, in the taxonomy of 21 risks of language models provided by Weidinger et al. (2022), the first category is "Discrimination, Hate speech and Exclusion". Given, the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al. (2018); Rudinger et al. (2018). It may be almost impossible to automatically filter a dataset or debias a model to be 100% free of unfair, bias or toxic content but the research community and other stakeholders may need to determine what levels are acceptable and if it should be a requirement to have human-in-the-loop methods. In this section, we discuss further methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies. We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free, as much as possible, from bias.
@@ 5.1 Methods to audit, measure, and evaluate fairness and bias
Embedding Association Tests (EATs)have been used in several studies Caliskan et al. (2017); Kurpicz-Briki (2020); Wolfe et al. (2023), starting with the introduction of Word Embedding Association Test (WEAT)by Caliskan et al. (2017) and improvements in Sentence Embedding Association Test (SEAT)May et al. (2019) and Relational Inner Product Association (RIPA)Ethayarajh et al. (2019). Despite its widespread use, WEAT has the disadvantage that it may systematically overestimate the bias in a model. Another embedding evaluation method is cosine similarity. It was used for zero-shot classification by Radford et al. (2021). It may also be used to audit fairness and bias by evaluating the similarity in image and text embeddings Wolfe et al. (2023). The visual tool, Gradient-weighted Class Activation Mapping (GRAD-CAM), generates a saliency map, which shows the most relevant regions of an image for given attributes Selvaraju et al. (2017); Wolfe et al. (2023). In an evaluation carried out by Wolfe et al. (2023), they discovered that the computed average saliency maps included only face regions for non-objectified images but both face and chest regions for objectified images, in a possible indication of sexual objectification bias. Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al. (2021); Birhane et al. (2024b)
@@
A recent metric introduced by Alkhaled et al. (2023) is bipol. It uses a two-step procedure in estimating bias in data Adewumi et al. (2023b); Pagliai et al. (2024). Bipol has the weakness that if the bias classifier is not accurate enough, false positives will weaken the evaluation score. Another measure is Area Under the Curve (AUC), as used by Meng et al. (2022) in the investigation of algorithmic fairness of mortality prediction, where they noted that Machine Learning (ML)methods obtain lower scores, usually, when it involves groups with higher mortality rates. Teo et al. (2024) proposed CLassifier Error-Aware Measurement (CLEAM), a framework for better performance in bias estimation, while Booth et al. (2021) measured gender bias using accuracy of Spearman rank-based correlation ($\rho$). Furthermore, Nozza et al. (2021) introduced the score “HONEST”, which was tested with respect to gender bias in text generation in 6 languages: Italian, French, Portuguese, Romanian, Spanish and English. It measures the probability that a language model will output hurtful text given a certain template and lexicon.
5.2 Datasets for Bias Evaluation
@@
Different datasets have been introduced for bias evaluation. Esiobu et al. (2023) introduced 2 novel datasets AdvPromptSet and HolisticBiasR, with which they evaluated 12 demographic dimensions for different LLM s. Ruzzante et al. (2022) introduced Sexual OBjectification and EMotion Database (SOBEM)for sexual objectification bias studies. It consists of 280 pictures of objectified and non-objectified female models with 3 different emotions and a neutral face. Bias Benchmark for QA (BBQ)is a question-sets dataset that employs templates crafted to reflect specific biases identified in society. It was introduced by Parrish et al. (2022) and aims to expose implicit prejudices that may exist against individuals from legally protected categories. BEAVERTAILS was introduced by Ji et al. (2024). It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exclusive. Discrimination, Stereotype, Injustice make up the second category in the list. Additional datasets for bias evaluation include RedditBias by Barikeri et al. (2021), RealToxicityPrompts, which comprises of 100K English sentences Gehman et al. (2020), HarmfulQ for zero-shot Chain of Thought (CoT)across stereotype benchmarks and harmful questions Shaikh et al. (2023), and Bold Dhamala et al. (2021).
5.3 Debiasing strategies on data, algorithms, procedures
@@ 5.3.1 Curate over Crawl
One important method to address bias in datasets will be to curate over crawl. This is especially so because web crawling has been the popular approach to getting internet data in the shortest possible time Birhane et al. (2021). The assumption of scale beats noise is the rationale for this approach by some researchers Jia et al. (2021). Unfortunately, this misconception about scaling does not only scale the quality part of the dataset but the noise along with it, no matter how small, as shown by Birhane et al. (2024b) when they observed 12% increase in hate content due to scaling. Clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
@@ 5.3.2 Counterfactual Data Augmentation (CDA)
Zhao et al. (2018) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods. It involves generating alternate examples of what exists (counterfactual) of data points to counter or mitigate bias. This method has gained attention in the field Meade et al. (2022); Barikeri et al. (2021).

### v2

@@ 4.2 LLM s
The relevant datasets, models, and challenges of bias affecting LLM s, as discussed in the literature, are summarized in Tables 3 and 4. The introduction of the Generative Pretrained Transformer (GPT-2) Radford et al. (2019) was a turning point in the language model landscape with its 1.5B parameters. As pointed out earlier, training such a large model required a lot of data and the Internet-sourced WebText was used for this purpose. Updated versions of the dataset have also been used for its recent successors, including GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023). Unfortunately, the attendant problems of bias witnessed in GPT-2 Gehman et al. (2020) have followed successive versions. This is also the case with other models, as expressed in Table 4, for the reasons that the datasets used in pretraining, some of which are given in Table 3, are largely from Internet sources. Furthermore, the architectures of the models, many of which are based on the Transformer architecture Vaswani et al. (2017), share similarity.
@@
In other works, Schramowski et al. (2022) show that moral directions, i.e. what is morally right or wrong to do, are present in LLM s. It is, however, highly debatable if models can be considered moral agents. In a collaborative effort, bench authors (2023) probed LLM s in the BIG-Bench of 200 tasks, including many that are related to bias. Also, Santurkar et al. (2023) explored whose opinions LLM s reflect while Harrer (2023); Nashwan and Abujaber (2023) investigated bias in LLM s within healthcare.
4.3 Mitigation Categories
Quantitative bias measurement and mitigation in NLP may be placed into 3 categories: preprocessing, as applicable in general ML Hort et al. (2024), intrinsic Caliskan et al. (2017); May et al. (2019), and extrinsic, as observed by Ramesh et al. (2023). The first, second and third involve quantifying and mitigating bias in the training data, in the trained model’s representation, and in the outputs of the downstream task of the model, respectively. More work has focused on the latter two than the first and gender bias than other dimensions Ramesh et al. (2023). For example, Delobelle et al. (2022) and Welbl et al. (2021) measured bias in pre-trained language models. One main reason why more work has been on the latter, according to Sun et al. (2019), is that debiasing an existing model to adjust the outputs usually only requires ‘patching’ the model instead of retraining with modified (debiased) or new data, which is usually costly.
Preuse
This may be considered the first step of the NLP preprocessing method for debiasing. It involves quantifying the amount of bias (or toxicity) in a given dataset by using a model, after the model has been trained on a different dedicated dataset, without necessarily attempting to mitigate the bias in the given dataset. It can be defined by Equation 1, where $B(d)$ is a bias metric that takes data as input and returns a scalar, $s$, as the score. Examples of works involving this are Alkhaled et al. (2023); Adewumi et al. (2023b). HateBERT Caselli et al. (2021) has been used in similar settings. Equation 1 is similar to that in Brunet et al. (2019) of differential bias, where they approximate the effect of removing small parts of training data on bias. Gender bias is one example of the kind of bias that can be estimated, as it has been observed that there are more male than female terms in many NLP datasets Sun et al. (2019); Alkhaled et al. (2023); Pagliai et al. (2024). The ability to first estimate quantitatively the bias in a given dataset provides the basis to be able to determine the level of success of mitigation.
$\mathit{B{\left(d\right)}}={s}$ (1)
4.4 Bias in multilingual AI:
@@
Some of the multimodal data in Table 3 contain multilingual data. These result in multilingual models and embeddings. For example, CC3M-35L, COCO-35L, and WebLI, which was used to train PaLI. WebLI is a mix of pre-training tasks with texts in 109 languages Chen et al. (2022). Some of these languages fall in the category of low-resource languages Adewumi et al. (2023a). Kurpicz-Briki (2020) reports statistically significant bias in German word embeddings based on the origin of a name in relation to pleasant and unpleasant words using WEAT Caliskan et al. (2017). In the study by Wambsganss et al. (2022), they found that the pretrained German language models, GermanBERT, GermanT5, and German GPT-2, had substantial conceptual, racial, and gender bias. This was also confirmed by Kraft et al. (2022), who observed sexist stereotypes in some of the models (e.g. family- and care-related terms were associated with female while crime and perpetrators were associated with male). Similarly, for Dutch, Delobelle et al. (2020) investigated gender and occupation biases in RobBERT (a Dutch RoBERTa Liu et al. (2019)), through a template-based association test Kurita et al. (2019); May et al. (2019). Huang et al. (2020b) also identified biases related to people’s origin and age in Italian, English, Polish, Portuguese, and Spanish, using a Twitter corpus.
@@ 5 Discussion
Although there are many limitations or risks of multimodal AI or LLM s Acosta et al. (2022); Adewumi et al. (2024b); Pettersson et al. (2024); Adewumi et al. (2024a), perhaps the issue of fairness and bias rank among the topmost Mehrabi et al. (2021). In addition, in the taxonomy of 21 risks of language models provided by Weidinger et al. (2022), the first category is ‘Discrimination, Hate speech and Exclusion’. Given the recurring challenges in this regard, we are of the view that existing tools for evaluating or handling fairness and bias in these systems need to be improved Zhao et al. (2018a); Rudinger et al. (2018). It may be almost impossible to automatically filter a dataset or debias a model to be 100% free of unfair, bias or toxic content but the research community and other stakeholders may need to determine what levels are acceptable and if it should be a requirement to have human-in-the-loop methods. In this section, we discuss methods to audit or evaluate fairness and bias, datasets for such evaluation, and debiasing strategies. We hope that such discussion will spur more researchers and stakeholders to see the critical importance of AI that is fair and free from bias, as much as possible.
@@ 5.1 Methods to audit, measure, and evaluate fairness and bias
Caliskan et al. (2017) introduced the Word Embedding Association Test (WEAT), which is based on the Implicit Association Test (IAT)Nosek et al. (2002). The IAT was designed to measure attitudes towards social groups. It showed implicit preference for White and young people over Black and old people, respectively. Furthermore, it showed the association of male terms with science while female terms were with family and arts. Embedding Association Tests (EATs)have been used in several studies Kurpicz-Briki (2020); Wolfe et al. (2023) and adapted with improvements in Sentence Embedding Association Test (SEAT)May et al. (2019) and Relational Inner Product Association (RIPA)Ethayarajh et al. (2019). Despite its widespread use, WEAT has the disadvantage that it may systematically overestimate the bias in a model. In vision models, Mandal et al. (2023b) used WEAT to audit CLIP by detecting and quantifying bias. Also, along the lines of the WEAT, Dev and Phillips (2019) introduced the Embedding Coherence Test (ECT)and Embedding Quality Test (EQT)and proposed methods for eliminating explicit bias. However, their method has the weakness that it is not able to remove implicit bias Friedrich et al. (2021).
@@
Another embedding evaluation method is cosine similarity. It was used for zero-shot classification by Radford et al. (2021). It may also be used to audit fairness and bias by evaluating the similarity in image and text embeddings Wolfe et al. (2023). The visual tool, Gradient-weighted Class Activation Mapping (GRAD-CAM), generates a saliency map, which shows the most relevant regions of an image for given attributes Selvaraju et al. (2017); Wolfe et al. (2023). In an evaluation carried out by Wolfe et al. (2023), they discovered that the computed average saliency maps included only face regions for non-objectified images but both face and chest regions for objectified images, in a possible indication of sexual objectification bias. The tool Not Safe For Work (NSFW)detector uses a tag alongside each image for filtering undesirable content Schuhmann et al. (2021); Birhane et al. (2024b)
@@
A recent metric introduced by Alkhaled et al. (2023) is bipol. It uses a two-step procedure in estimating bias in data Adewumi et al. (2023b); Pagliai et al. (2024). Bipol has the weakness that if the bias classifier is not accurate enough, false positives will weaken the evaluation score. Another measure is Area Under the Curve (AUC), as used by Meng et al. (2022a) in the investigation of algorithmic fairness of mortality prediction, where they noted that ML methods obtain lower scores, usually, when it involves groups with higher mortality rates. Teo et al. (2024) proposed CLassifier Error-Aware Measurement (CLEAM), a framework for better performance in bias estimation, while Booth et al. (2021) measured gender bias using accuracy of Spearman rank-based correlation ($\rho$). Furthermore, Nozza et al. (2021) introduced the score “HONEST”, which was tested with respect to gender bias in text generation in 6 languages: Italian, French, Portuguese, Romanian, Spanish and English. It measures the probability that a language model will output hurtful text given a certain template and lexicon.
Datasets for Bias Evaluation
@@
Different datasets have been introduced for bias evaluation. Liang et al. (2021a) introduced MultiBench, a unified multimodal benchmark that spans 15 datasets, 10 modalities, and 20 prediction tasks. FairFace was introduced by Karkkainen and Joo (2021). The dataset was designed to mitigate racial bias in multimodal AI, collected from the Yahoo Flickr Creative Commons 100 Million (YFCC100M)dataset Thomee et al. (2016) and contains 108,501 images, balanced across the following races: Black, White, Indian, Southeast Asian, East Asian, Middle Eastern, and Latino. Esiobu et al. (2023) introduced 2 novel datasets AdvPromptSet and HolisticBiasR, with which they evaluated 12 demographic dimensions for different LLM s. Ruzzante et al. (2022) introduced Sexual OBjectification and EMotion Database (SOBEM)for sexual objectification bias studies. It consists of 280 pictures of objectified and non-objectified female models with 3 different emotions and a neutral face. Bias Benchmark for QA (BBQ)is a question-set dataset that employs templates crafted to reflect specific biases identified in society. It was introduced by Parrish et al. (2022) and aims to expose implicit prejudices that may exist against individuals from legally protected categories.
BEAVERTAILS was introduced by Ji et al. (2024). It assesses question-answer pairs, tested on LLM s, with regards to 14 different harm categories, which are not exhaustive. ‘Discrimination, Stereotype, Injustice’ makes up the second category in the list. Additional datasets for bias evaluation include RedditBias by Barikeri et al. (2021), RealToxicityPrompts, which comprises of 100K English sentences Gehman et al. (2020), HarmfulQ for zero-shot Chain of Thought (CoT)across stereotype benchmarks and harmful questions Shaikh et al. (2023), and BOLD Dhamala et al. (2021). Porgali et al. (2023) also introduced the Casual Conversations dataset (version 2) containing 26,467 videos of 5,567 unique participants from 7 different countries, representing a wide range of demographics, for bias and robustness evaluation of LMM s that are vision and audio models.
5.2 Debiasing strategies
Although there’s no silver bullet to solving the challenges of fairness and bias in the data and models of multimodal AI, we believe a combination of two or more of the following strategies on the relevant datasets or models in Tables 3 and 4 will go a long way in mitigating bias in AI generally.
@@ 5.2.1 Curate over Crawl
One important method to address bias in datasets will be to curate rather than crawl. This is especially so because web crawling has been the popular approach to getting Internet data in the shortest possible time Birhane et al. (2021). The assumption of scale beats noise is the rationale for this approach by some researchers Jia et al. (2021). Unfortunately, this misconception about scaling does not only scale the quality part of the dataset but the noise along with it, no matter how small, as shown by Birhane et al. (2024b) when they observed 12% increase in hate content due to scaling. On the other hand, clearly, despite the advantage of curation, one hurdle to overcome with the curate over crawl approach will be the issue of scaling.
In addition, for better quality data collection, researchers like Jo and Gebru (2020) advocate that AI practitioners should build on the practices of those in the field of archives and libraries and have a public mission statement to guide their data collection practice Weinberg (2022). Furthermore, Gebru et al. (2021) encourage, through ‘datasheets for datasets’, standardized processes for documenting important aspects of the dataset creation, including motivation, composition, funding, collection and use cases, with the potential to mitigate unwanted biases, though this approach has its limitations because individuals included or affected by the datasets are not necessarily empowered to influence them Weinberg (2022).
@@ 5.2.2 Counterfactual Data Augmentation (CDA)
Zhao et al. (2018a) used CDA to show that it removes bias with minimal performance degradation on coreference benchmarks when combined with existing word-embedding debiasing methods. It involves generating alternate examples of what exists (counterfactual) of data points to counter or mitigate bias. This method has gained attention in the field Meade et al. (2022); Barikeri et al. (2021). It is sometimes called ‘gender-swapping’ in the specific case of gender bias Sun et al. (2019). CDA has been shown to be effective in tackling bias in coreference resolution, as mentioned earlier, as it reduced the difference in F1 scores between pro-stereotypical and antistereotypical evaluations Sun et al. (2019). Dev and Phillips (2019) also embraced this neutralizing approach in their work by flipping gendered terms with their counterparts. For example, a sentence like ‘he was a doctor’ would be flipped to ‘she was a doctor’. Despite the advantages, as pointed out earlier about the less-discussed preprocessing bias mitigation method, completely debiasing textual data can be complicated and intricate. Some of the challenges are that the size of the data increases significantly, thereby increasing model training time, and naive term-swapping can create more challenges of unrealistic or nonsensical samples in the data, e.g. ‘she has hot flashes because of menopause’ to ‘he has hot flashes because of menopause’ Sun et al. (2019).
@@ 5.2.3 Improved Filtering
5.2.4 Linear Projection
This method projects all words $w\in W$ orthogonally to the bias vector, ensuring the updated set has no component along the bias vector, $v_{B}$, as given in Equation 2 Bolukbasi et al. (2016); Dev and Phillips (2019).
$\mathit{w\textsuperscript{´}}=\mathit{w-\pi_{B}{\left(w\right)}}=\mathit{w-\langle w,v_{B}\rangle v_{B}}$ (2)
The span that results becomes less by 1 in the total dimensions, say from 300 to 299, which will have negligible effects on the generalizability of the embeddings. As an example with gender bias, subtraction with linear projection of gender terms from embeddings will make them close. Gendered word-pairs that have few word sense (e.g. he - she) and (him - her) can have close enough identifcal positions in the vector space after debiasing.
5.2.5 Debiasing Word Embeddings
Removing gender bias from word embeddings can take one of two approaches: (1) removing gender subspace Bolukbasi et al. (2016) and (2) learning gender-neutral embeddings Zhao et al. (2018b). The two approaches may not be adequate for embeddings that are not based on Euclidean space since cosine similarity will no longer apply Sun et al. (2019). The first approach modifies an embedding based on the combined properties of word embeddings that gender bias can be captured by a direction and neutral words are linearly separable from gendered words Bolukbasi et al. (2016) while the second approach preserves gender information in some dimensions but compels other dimensions of the word embedding to be free of such Zhao et al. (2018b).
5.2.6 Adapters
A post-processing method for bias mitigation that is based on Adapter modules Houlsby et al. (2019), called Debiasing with Adapter Modules (DAM), was introduced by Kumar et al. (2023). They encapsulate different bias mitigation functionalities and can be integrated when desired in a model, similarly to how AdapterFusion Pfeiffer et al. (2021) is carried out in multi-task learning. DAM trains the main adapter and the bias mititgation adapters independently before combining them. DAM follows an earlier adapter-based debiasing method, called Adapter-based DEbiasing of LanguagE Models (ADELE), performed in the work by Lauscher et al. (2021). Their approach involved the additional use of CDA.
5.2.7 Additive Residuals
To address the skewed distribution of different identity groups in the training data used in LMM s, Seth et al. (2023) introduced Debiasing with Additive Residuals (DeAR)to learn additive residual image representations. This minimizes the representations’ capacity to distinguish among different identity groups, thereby offering fairer output.
5.2.8 Continued Pretraining
Fatemi et al. (2023) built on the continued pretraining concept, which is sometimes used for gender bias mitigation with a small gender-neutral dataset de Vassimon Manela et al. (2021), by introducing GEnder Equality Prompt (GEEP)such that it reduces catastrophic forgetting, which is a likely event in continued pretraining Kirkpatrick et al. (2017). GEEP achieves this by freezing the entire model before updating the embeddings. Furthermore, Cabello et al. (2023) showed that continued pretraining on gender-neutral data improves fairness by reducing group disparities in some language-vision tasks.
5.2.9 Adversarial Learning
Yan et al. (2020) used adversarial learning for bias mitigation as proposed by Zhang et al. (2018). They added a discriminator to jointly learn with the predictor for the sensitive attributes. In this approach, the generator prevents the discriminator from identifying gender in a task.
5.2.10 Gender Tagging
In Machine Translation (MT), gender-tagging may be used Vanmassenhove et al. (2018). It involves the addition of gender tags to the beginning of data samples to identify the gender of the source data, e.g. ‘FEMALE I’m travelling tomorrow’. Apparently, for more complex sentences this approach may become more challenging Sun et al. (2019).
5.2.11 FairDistillation
To address bias across languages, Delobelle and Berendt (2022) introduced FairDistillation. It is a cross-lingual method that is based on knowledge distillation Hinton et al. (2015) by creating smaller language models from large ones to control for stereotypical and representational biases.
5.2.12 DEBIE
Friedrich et al. (2021) introduced DEBIE as a platform for measuring and mitigating implicit and explicit bias in word embeddings. The mitigation methods are more specific to NLP and not available in general purpose library such as AI Fairness 360 (AIF360)Bellamy et al. (2019). DEBIE is a collection of commonly used bias data tools and word embeddings, including fastText Bojanowski et al. (2017), GloVe Pennington et al. (2014), Continuous Bag-of-Words (CBoW)Mikolov et al. (2013), and the WEAT test.